OpenStudy (anonymous):

decompose 1/x^2(x^2+1)^2

OpenStudy (unklerhaukus):

into partial fractions?

OpenStudy (anonymous):

yes please :(

OpenStudy (unklerhaukus):

The first term in the denominator is the x^2 term

OpenStudy (anonymous):

ohhhh that's why i can't get the right answer.. TY very much!!!

OpenStudy (unklerhaukus):

what partial fraction terms will the x^2 term decompose into?>

OpenStudy (anonymous):

wait so i will write the first term as A/x^2?

OpenStudy (perl):

Using your partial fraction table. it might help to write x^2 as (1*x+0)^2

OpenStudy (unklerhaukus):

for should two fractions from the first term

OpenStudy (anonymous):

so i will write it as A/x + B/x^2 for the first term?

OpenStudy (unklerhaukus):

yes, that's good. now lets look at the second term

OpenStudy (anonymous):

so for the second term it will be (Cx+D)/(x^2+1) + (Ex+F)/(x^2+1)^2?

OpenStudy (unklerhaukus):

Yes, great work, now we just have to find A, B, C, D, E, & F

OpenStudy (anonymous):

Ok i'll be back later

OpenStudy (unklerhaukus):

So far, we have \[\frac1{x^2(x^2+1)^2}=\frac Ax + \frac B{x^2}+\frac{Cx+D}{x^2+1} + \frac{Ex+F}{(x^2+1)^2}\]

OpenStudy (anonymous):

okay okay...i'm already solving for the values.

OpenStudy (anonymous):

is this right? A=0, B=1, C=-1, D=0, E=0, and F=1?